A student stands at the edge of a cliff and throws a stone horizontally over the edge with a speed of 18m/s.  The cliff is 50.m above a flat horizontal beach.  

How long after being released does the stone strike the beach below the cliff?  

With what velocity does it land?



1
Expert's answer
2021-05-11T16:35:00-0400

(a) We can find the time that the stone takes to reach the beach below the cliff from the kinematic equation:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}},""t=\sqrt{\dfrac{2\cdot50\ m}{9.8\ \dfrac{m}{s^2}}}=3.19\ s."

(b) Let's first find the vertical component of the stone's velocity:


"v_y=-gt=-9.8\cdot3.19\ s=-31.3\ \dfrac{m}{s}."

The sign minus means that the vertical velocity of the stone directed downward.

Finally, we can find with what velocity the stone lands from the Pythagorean theorem:


"v=\sqrt{v_x^2+v_y^2},""v=\sqrt{(18.0\ \dfrac{m}{s})^2+(-31.3\ \dfrac{m}{s})^2}=36.1\ \dfrac{m}{s}."